OpenStudy (anonymous):

Solve x^2 + 4x + 8 = 0

OpenStudy (anonymous):

x^2 +4x = -8 x(x+4) = -8 x = -8 and x = -12

OpenStudy (anonymous):

x1={(-4)+sqrt(-48)}/2 , x2={(-4)-sqrt(-48)}/2 x1=-2+i(2sqrt(3)); x2=-2-i(2sqrt(3))